OpenStudy (anonymous):

-2x+3y+z=-1 -x+7y-5z=-6 -3x+5y+z=-2 whats z,x, and y?

OpenStudy (anonymous):

You can choose y arbitrarily and set x = 1+2*y, z = 1+y